# Bayesian Concept Bottleneck Models with LLM Priors

### v1

@@ 3.2 Bayesian variable selection with LLMs
In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) (Liu et al.,, 2000). Here, the LLM proposes a batch of candidate candidates, and then we sample a candidate from this batch based on their posterior probabilities and consider the sampled candidate for acceptance. This is more cost-effective as it lets us batch concept annotations, as described later.
@@
1:function Multi-SS-MH-Update($\vec{c}$, $k$)
@@
Assuming that the LLM’s partial posterior across all iterations is consistent with some prior distribution $p(\vec{C})$, we can show that the Markov chain defined by this procedure has the posterior $p(\vec{c}|\mathbf{y},\mathbf{X})$ as its stationary distribution (see Proposition I.1 in the Appendix). More importantly, we can prove that the procedure still converges to the true concepts even in the likely scenario where there is no valid prior.
@@
Suppose the data is IID. Let $L(\vec{c})\coloneqq\max_{\vec{\theta}}\mathbb{E}\{\log p(Y|X,\vec{\theta},\vec{c})\}$ and $\mathcal{C}^{*}\coloneqq\operatorname{argmax}_{\vec{c}}L(\vec{c}).$ For sample size $n$, let $\Pi_{n}$ denote the set of stationary distributions of the Markov chain defined by running Algorithm 1 with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update. Assuming a Gaussian prior for $\vec{\theta}$ and under regularity conditions (see Assumption H.1 in the Supplement), then $\inf\{\pi(\mathcal{C}^{*})\colon\pi\in\Pi_{n}\}\to 1$ in probability as $n\to\infty$.
3.3 BC-LLM
We now describe the main steps in BC-LLM (Fig 1): the procedure is initialized in Step 0, and each iteration (Steps 1 to 4) samples a CBM. The samples then describe the posterior distribution over CBMs and form an ensemble prediction model (after removing those from the burn-in period). Depending on one’s use case, the search over concepts can be guided by the LLM’s prior knowledge about the target $Y$ by injecting this information into the prompts (e.g. “we are trying to predict readmission…”) or in an “unbiased” fashion by masking the meaning of $Y$ (e.g. “we are trying to predict some label $Y$…”). Example prompts and other implementation details are provided in the Appendix.
Step 0: Initialization and keyphrase extraction
@@ 3.3 BC-LLM
We first use an LLM to extract “keyphrases” describing each observation $x$, which will later be used by the LLM to brainstorm candidate concepts. Here, we use “keyphrases” to refer to short phrases that describe the observation: keyphrases for images describe what’s in the image (e.g. blue eyes), while keyphrases for text data can be phrases that appear in the text or summarize the main ideas (e.g. diabetes) (Papagiannopoulou and Tsoumakas,, 2020). Unlike concepts, keyphrases do not define a function that outputs a value, though they can help one come up with a concept (e.g. “diabetes” is a keyphrase whereas the binary questions “Does the patient currently have diabetes?” or “Does the patient have a family history of diabetes?” are concepts). In addition, this step does not restrict the set of keyphrases one can extract to a prespecified list, which allows for a wide exploration of keyphrases. The tradeoff is that the extracted keyphrases may be, and likely are, an incomplete summary of each observation.
To initialize the CBM sampling procedure, these keyphrases are converted into numeric vectors to fit a (penalized) LR model for the target $Y$, similar to the process for fitting a bag-of-words model. An LLM analyzes the top keyphrases in this LR model to generate an initial set $\vec{c}$ of $K$ concepts. The LLM also extracts these concepts from each observation.
Step 1: Drop a concept & fit a keyphrase model
Each iteration proposes candidate concepts based on some data subset $S$ to replace the $k$th concept for $k\in\{1,\cdots,K\}$. Given the context limits of LLMs, it is unlikely that an LLM can generate a candidate concept from the partial posterior distribution by reviewing each observation in $S$ individually. As such, we will generate a summary by highlighting the top keyphrases in $S$ for predicting $Y$, on top of the existing $K-1$ concepts $\vec{c}_{-k}$.
For continuous outcomes, this can be accomplished by finding the top keyphrases associated with the residuals of a model that predicts $Y$ given $\vec{c}_{-k}$. As the residuals are noisy for binary/categorical outcomes, we fit a (multinomial) LR model of $Y$ instead, where the input features are the keyphrases (again encoded using a bag-of-words approach) and annotated concepts $\vec{c}_{-k}$. The parameters of this “keyphrase model” $f$ are the coefficients for the keyphrases and concepts, denoted $\beta_{W}$ and $\beta_{\vec{c}_{-k}}$, respectively. Given the high-dimensionality of the keyphrases, we penalize $\beta_{W}$ using some penalty function $J$ and estimate $f$ by solving
$\min_{\beta_{\vec{c}_{-k}},\beta_{W}}\frac{1}{n}\sum_{i=1}^{n}\ell\left(y_{i},f\left(x_{i}|\beta_{\vec{c}_{-k}},\beta_{W}\right)\right)+\lambda J(\beta_{W}),$
where $\ell$ is the logistic loss and penalty parameter $\lambda\geq 0$ is selected by cross-validation. We found the ridge penalty (i.e., $J(\beta_{W})=\|\beta_{W}\|_{2}^{2}$) to work well in practice, though other choices for $J$ may be used instead.
@@
Step 2: Query the LLM for candidate concepts Given the summary of the top keyphrases generated from Step 1, the LLM is asked to propose $M$ candidates for the $k$-th concept from the partial posterior $p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$. In addition, we ask the LLM to provide the value of the partial posterior probability for each candidate concept. For instance, if the top keyphrases include “chronic kidney disease” and “acute kidney injury,” the LLM may propose a concept like “Does the patient have a history of kidney disease?” with an assigned partial posterior probability of 0.4.
@@
Step 3: Concept annotation Given the batch of candidate concepts, we use the LLM to extract their values for each observation $x$, i.e. $\phi_{c}(x)$ for $c\in\{\check{c}^{(1)}_{k},\ldots,\check{c}^{(M)}_{k}\}$. Because multiple-try Metropolis-within-Gibbs proposes a batch of candidate concepts, we can batch their extractions into a single LLM query for each observation. This makes BC-LLM more cost/query-efficient than ordinary Metropolis-within-Gibbs, which outputs only a single concept per iteration and thus cannot batch extractions.
@@
[…]
@@
To estimate the integral (3), one approach is to sample from the posterior on $\vec{\theta}$, but this can be computationally expensive since one would need to implement posterior sampling in inner and outer loops. Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation for (3) (Lewis and Raftery,, 1997), which is discussed in more detail in the Appendix. This essentially approximates the partial and full posterior distributions with Gaussian distributions. As the posterior distribution of LR parameters are asymptotically normal, this approximation becomes increasingly accurate as $n\rightarrow\infty$.
3.4 Computational cost
@@
The key computational cost of BC-LLM is from querying the LLM. Running BC-LLM requires $O(nTK)$ queries, where $T$ is the number of epochs. $K$ is typically small by design, because large $K$ leads to less interpretable CBMs. One may be concerned by $T$, as standard Bayesian inference procedures typically draw thousands of posterior samples. Nevertheless, combined with a greedy warm-start procedure, we found that a small $T$ to be quite effective; our experiments go as low as $T=5$. Generally speaking, if the goal is solely prediction accuracy, a small $T$ may suffice. If the goal is uncertainty quantification for relevant concepts, a larger $T$ (e.g., 10) is helpful for more complete exploration of concepts.
Figure 2: MIMIC evaluation: (a) Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts. Error bars indicate the 95% CI. (b) Dendrograms of concepts learned by BC-LLM with 100 and 800 training observations (left and right, respectively). Labels are shortened concept questions, which generally have the format “Does the note mention the patient having…?” Highlighted labels correspond to the true, color-coded concepts in Section 4.2. Numbers in parentheses describe the proportion of posterior samples with the concept.
4 Experimental results
4.1 Experimental setup
We compare BC-LLM against One-pass summarization, where an LLM performs post-hoc distillation of the keyphrase model into a CBM (Pham et al.,, 2023), and Boosting, where an LLM iteratively adds concepts to a CBM by analyzing observations with large residuals (Ludan et al.,, 2023; Liu et al., 2024a, ). We also compare it to black-box and semi-interpretable models to assess whether BC-LLM can close the performance gap. When the true concepts are known, we also fit an Oracle CBM.
Methods were evaluated with respect to predictive performance as measured by AUC and uncertainty quantification as measured by log likelihood (higher is better). When true concepts are known, we also quantify concept selection rates using “concept precision,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(\hat{C}_{k}\notin\vec{c}^{*}\mid\mathbf{y},\mathbf{X})$, and “concept recall,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(c_{k}^{*}\notin\hat{\vec{C}}\mid\mathbf{y},\mathbf{X})$, where $\hat{\vec{C}}$ is the posterior distribution over concept sets from BC-LLM or a single set of concepts learned by a non-Bayesian procedure. Quantifying a “correct” concept match requires going beyond exact string matching, as there are many possible refinements for a single concept. For instance, for the concept “Does the note mention the patient smoking in the present or the past?”, there are rephrasings (“Does this note mention the patient currently smoking or having smoked?”), polar opposites (“Does the note not mention the currently smoking or having previously smoked?”), or closely overlapping concepts (“Does the note mention the patient currently smokes?”). To quantify if a learned concept was correct, we used the following rule: if the absolute value of the Pearson correlation for LLM annotations of two concepts exceeds 50%, the two concepts are sufficiently similar. We manually confirmed that any pair of learned and true concepts that passed this rule were always closely related.
@@ 4.1 Experimental setup
Experiments used GPT-4o-mini (OpenAI,, 2023) and a protected health information (PHI)-compliant version of GPT-4o for real-world clinical notes in Section 4.4. Additional implementation details and experimental settings are in the Appendix.
4.2 Simulation study: Learning CBMs for clinical notes from MIMIC
@@
We first compared BC-LLM against existing methods in terms of predictive performance and concept selection rates. To assess how well the methods perform even when the LLM has no prior knowledge about the target, we simulated $Y$ using a LR model with five Social Determinants of Health as inputs, which were annotated from real-world patient notes in MIMIC-IV (Johnson et al.,, 2023; Ahsan et al.,, 2021):

### v2

@@ 3.2 The Split-sample update: Propose with the LLM’s partial posterior
1:function Multi-$S$-MH-Update($\vec{c}$, $k$)
@@
In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) [45]. Here, the LLM proposes a batch of candidates. We sample a candidate from this batch based on their posterior probabilities and consider the sampled candidate for acceptance. This is more cost-effective as it lets us batch concept annotations, as described later.
Assuming that the LLM’s partial posterior across all iterations is consistent with some prior distribution $p(\vec{C})$, we can prove that the Markov chain defined by this procedure has the posterior $p(\vec{c}|\mathbf{y},\mathbf{X})$ as its stationary distribution (see Proposition H.1 in the Appendix). Since LLMs are known to be imperfect Bayesian inference engines [39], this assumption is at best only approximately satisfied. Nevertheless, we can prove that the procedure still converges to the true concepts even when this assumption does not hold:
@@
Suppose the data is IID. Let $L(\vec{c})\coloneqq\max_{\vec{\theta}}\mathbb{E}\{\log p(Y|X,\vec{\theta},\vec{c})\}$ and $\mathcal{C}^{*}\coloneqq\operatorname{argmax}_{\vec{c}}L(\vec{c}).$ For sample size $n$, let $\Pi_{n}$ denote the set of stationary distributions of the Markov chain defined by running Algorithm 1 with SS-MH-Update or Multi-SS-MH-Update instead of MH-Update. Assuming a Gaussian prior for $\vec{\theta}$ and under regularity conditions (see Assumption G.1), then $\inf\{\pi(\mathcal{C}^{*})\colon\pi\in\Pi_{n}\}\to 1$ in probability as $n\to\infty$.
3.3 Putting it together: BC-LLM
We now put all the pieces together. Mathematically, BC-LLM runs Metropolis-within-Gibbs (Algorithm 1) with a split-sample update (Algorithm 2 or 3). We translate this into a step-by-step recipe (Fig 1): After initializing BC-LLM in Step 0, each iteration of Metropolis-within-Gibbs (Steps 1 to 4) performs a split-sample update to output a CBM. The CBMs form a posterior distribution, as well as an ensemble prediction model. Example prompts and implementation details are in the Appendix.
@@ 3.3 Putting it together: BC-LLM
Step 0: Initialization and keyphrase extraction We first use an LLM to extract “keyphrases” from each observation $x$, which will be used by the LLM to brainstorm candidate concepts. Here, “keyphrases” refer to short phrases: keyphrases for images describe what’s in the image (e.g. blue eyes), while keyphrases for text data can be direct quotes or summarizations (e.g. diabetes) [46]. Keyphrases can be thought of as rough concepts, as they do not define a function but can help an LLM come up with a formal concept question (e.g. “diabetes” is a keyphrase whereas “Does the patient currently have diabetes?” is a concept). This brainstorming phase does not restrict the set of keyphrases the LLM can extract to allow for a wide exploration of concepts. The tradeoff is that the extracted keyphrases may be an incomplete summary of each observation. To initialize the CBM sampling procedure, these keyphrases are used to fit a (penalized) LR model for the target $Y$, similar to fitting a bag-of-words (BoW) model. An LLM then analyzes the most predictive keyphrases to generate an initial set of $K$ concepts and extracts these concepts from each observation.
Step 1: Drop a concept & fit a keyphrase model Each iteration proposes candidate concepts based on some data subset $S$ to replace the $k$th concept for $k\in\{1,\cdots,K\}$. Given the context limits of LLMs, having an LLM review each observation in $S$ individually is ineffective. Instead, we generate a summary by highlighting the top keyphrases in $S$ for predicting $Y$, on top of the existing K1K-1 concepts $\vec{c}_{-k}$. For continuous outcomes, this can be accomplished by finding the top keyphrases associated with the residuals of a model that predicts $Y$ given $\vec{c}_{-k}$. For binary/categorical outcomes, we instead fit a (multinomial) LR model of $Y$, where the inputs are keyphrases encoded using BoW and annotated concepts $\vec{c}_{-k}$. Let the coefficients in this “keyphrase model” $f$ for keyphrases and concepts be denoted $\beta_{W}$ and $\beta_{\vec{c}_{-k}}$, respectively. We fit it by solving $\min_{\beta_{\vec{c}_{-k}},\beta_{W}}\frac{1}{n}\sum_{i=1}^{n}\ell\left(y_{i},f\left(x_{i}|\beta_{\vec{c}_{-k}},\beta_{W}\right)\right)+\lambda J(\beta_{W})$, where $\ell$ is the logistic loss, $\beta_{W}$ is penalized with some penalty function $J$, and penalty parameter $\lambda$ is selected by cross-validation. We found the ridge penalty to work well in practice but other penalties may be used instead.
@@
Step 2: Query the LLM for candidate concepts Given the summary of the top keyphrases generated from Step 1, the LLM is asked to propose $M$ candidates for the $k$-th concept. In addition, we ask the LLM to provide the value of the partial posterior probability for each candidate concept. For instance, if the top keyphrases include “chronic kidney disease” and “acute kidney injury,” the LLM may propose a concept like “Does the patient have a history of kidney disease?” with an assigned partial posterior probability of 0.4.
@@
Step 3: Concept annotation Given candidate concepts, we use the LLM to extract their values using a zero-shot approach, i.e. $\phi_{c}(x)$ for $c\in\{\check{c}^{(1)}_{k},\ldots,\check{c}^{(M)}_{k}\}$. Because multiple-try Metropolis-within-Gibbs proposes a batch of candidate concepts, we can batch their extractions into a single LLM query for each observation. This makes BC-LLM more cost/query-efficient than ordinary Metropolis-within-Gibbs, which outputs only a single concept per iteration and thus cannot be batched. Note that letting the LLM occasionally output probabilities when it is unsure can be helpful, though it is critical to avoid information leakage (e.g. no joint training of concept extraction and label prediction).
@@
$p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})=\int p(\mathbf{y}_{S^{c}}|\vec{\theta},\vec{c},\mathbf{X})p(\vec{\theta}|\mathbf{y}_{S},\vec{c},\mathbf{X})d\vec{\theta}.$ (1)
@@
To estimate the integral (1), one approach is to sample from the posterior on $\vec{\theta}$, but this can be computationally expensive since one would need to implement posterior sampling in inner and outer loops. Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation for (1) [47], which is discussed in more detail in the Appendix. As the posterior distribution of LR parameters are asymptotically normal, this approximation becomes increasingly accurate as $n\rightarrow\infty$.
@@
Computational cost. BC-LLM performs $O(nTK)$ LLM queries, where $T$ is the number of outer loops. $K$ is typically chosen to be small, because a large $K$ yields less interpretable CBMs. Although standard Bayesian procedures typically draw thousands of posterior samples, we found that a small $T$ was quite effective in practice when combined with a greedy warm-start procedure; our experiments all use $T=5$. Generally speaking, a small $T$ suffices for achieving high prediction accuracy, a larger $T$ (e.g., 10) is helpful for quantifying the uncertainty of relevant concepts. Critically, running BC-LLM is significantly more cost-effective than standard CBMs, which require $O(nW)$ queries where $W$ is the number of pre-specified concepts and must be large (typically hundreds or thousands) to avoid missing relevant concepts [8]. In our experiments, BC-LLM was quick to run, as one iteration takes less than a minute for text data and a few minutes for image data on a normal laptop.
4 Experiments
@@
We evaluated BC-LLM in across three domains and modalities: classifying birds in images (Section 4.1), simulated outcomes from clinical notes (Section 4.2), and readmission risk in real-world clinical data (Section 4.3). The experiments below used GPT-4o-mini [48] and Section 4.3 used a protected health information-compliant version of GPT-4o for real-world clinical notes. The Appendix includes results using other LLMs, implementation details, and experimental settings.
Baselines. We compared against CBMs fit using the standard approach where humans select and annotate concepts (Human+CBM) [5], CBMs where an LLM brainstorms concepts after seeing which keyphrases are most associated with the label but without any iteration (LLM+CBM) [49, 10], and CBMs fit using a boosting algorithm that iteratively adds concepts by having an LLM analyze observations with large residuals (Boosting LLM+CBM) [17, 24]. We include black-box and semi-interpretable models to assess the tradeoff between interpretability and accuracy. For the bird-classification dataset, we also compare to a CBM that uses LLM-suggested concepts rather than keyword-based concepts; we use the 370 LLM-suggested concepts from [8] and extract the values of these concepts with an LLM (LLM+CBM (No keyphrases)).
Evaluation metrics. Methods were evaluated with respect to predictive performance as measured by AUC and/or accuracy and uncertainty quantification as measured by Brier score. When true concepts are known, we also quantify concept selection rates using “concept precision,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(\hat{C}_{k}\in\vec{c}^{*}\mid\mathbf{y},\mathbf{X})$, and “concept recall,” as defined by $\frac{1}{K}\sum_{k=1}^{K}p(c_{k}^{*}\in\hat{\vec{C}}\mid\mathbf{y},\mathbf{X})$, where $\hat{\vec{C}}$ is the posterior distribution over concept sets from BC-LLM or a single set of concepts learned by a non-Bayesian procedure. Concept matches were verified manually (see Appendix for details).
4.1 CBMs for classifying bird images
We first evaluate how well BC-LLM can learn concepts that differentiate between bird species within the same family in the standard CUB-birds image dataset [50]. Grouping the 200 bird species into their respective families, this resulted in 37 prediction tasks. We considered family-based prediction tasks because there likely exists a parsimonious CBM within each bird family, in contrast to the task of predicting between all 200 species. To assess how well BC-LLM performs in settings with limited prior knowledge, we do not tell the LLM that we are predicting bird species. That is, we replace the labels with one-hot encodings and simply tell the LLM to find concepts for predicting $Y$. Consequently, the LLM must search over a larger space of concepts.
The data is split 50/50 between training and testing. The number of concepts $K$ learned for each task was set to the number of classes, but no smaller than 4 and no greater than 10. For the black-box comparator, we fine-tuned the last layer of ResNet50 pre-trained on ImageNetV2 [51, 52]. Human+CBM was trained on the 312 human-annotated features available in the CUB-birds dataset.
Table 1: Performance of CBMs and black-box models (ResNet) for classifying bird species.
In-distribution OOD
Method Accuracy ($\uparrow$) AUC ($\uparrow$) Brier ($\downarrow$) Entropy ($\uparrow$)
BC-LLM 0.680 (0.614, 0.747) 0.874 (0.840, 0.907) 0.428 (0.357, 0.500) 0.865 (0.693, 1.036)
LLM+CBM 0.640 (0.573, 0.707) 0.810 (0.768, 0.853) 0.452 (0.377, 0.528) 0.663 (0.474, 0.852)
Boosting LLM+CBM 0.538 (0.463, 0.614) 0.722 (0.673, 0.772) 0.577 (0.499, 0.654) 0.842 (0.630, 1.054)
Human+CBM 0.658 (0.591, 0.725) 0.835 (0.791, 0.879) 0.499 (0.414, 0.584) 0.758 (0.558, 0.959)
LLM+CBM (No keyphrases) 0.555 (0.488, 0.623) 0.759 (0.713, 0.805) 0.651 (0.548, 0.754) 0.626 (0.495, 0.757)
ResNet 0.664 (0.613, 0.716) 0.853 (0.821, 0.885) 0.457 (0.398, 0.516) 0.914 (0.748, 1.079)
BC-LLM achieved higher accuracy and better calibration on average, compared to all the other fully interpretable CBMs (Table 1). Notably, soft CBMs no longer performed as well once the concepts were constrained to be fully interpretable. BC-LLM also outperformed CBMs trained on human collated and annotated features, potentially because it was able to consider additional concepts. More critically, BC-LLM outperformed ResNet50 in both accuracy and calibration, as ResNet50 easily overtrained on the small bird datasets. In contrast, BC-LLM could identify relevant concepts for a simple parametric model, leading to faster model convergence.
We also assessed BC-LLM’s robustness in OOD settings, as the Bayesian framing should improve uncertainty quantification (Table 1). For each model, we used it to classify birds for a species it was not trained on. Compared to other CBM methods, BC-LLM was more appropriately unsure, yielding higher entropy across the output classes. We note that ResNet model had the highest entropy, though it has the advantage of not being limited to extracting a small set of interpretable concepts. To illustrate the benefit of using fully interpretable CBMs in OOD settings, Figure 2 (right) shows the difference when applying the Bunting bird CBM to an actual bunting bird versus an OOD sample. Whereas the CBM extracted zeros and ones for the actual bunting bird, it was unsure how to extract the concept “Does this image show vibrant feathers?” for the image of a dog wearing rainbow wings. The LLM thus outputted a probability of 0.5, citing that “the image contains vibrant colors but the wings do not actually contain feathers.” This shows how having an LLM output its reasoning when answering concrete concept questions can reveal situations that need human intervention.
Finally, we visualize the posterior distribution over concepts using hierarchical clustering (Fig 2 left). Continuing with the Bunting bird example, we see that BC-LLM converges towards distinguishing concepts as training data increases and away from vague or non-specific concepts (e.g. “sharp beaks” apply to all bunting birds). Moreover, the number of unique concepts decreases with the amount of training data, because it has converged to concepts that truly distinguish the different species.
Figure 2: Example of BC-LLM classifying Bunting birds. (Left) Left and right dendrograms are learned concepts when given 1/3 versus 3/3 of the training data, respectively. Labels are shortened concept questions, generally of the form “Does the image depict…?” Proportion of posterior samples with the concept are shown in parentheses. Highlighted labels are distinguishing bird features. (Right) Application of BC-LLM to an actual bunting bird (top) versus a dog pretending to be one (bottom).
4.2 CBMs for classifying clinical notes with simulated outcomes
Figure 3: MIMIC results: (Left) Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts with 95% CI. (Right) Dendrograms of concepts learned by BC-LLM with 100 and 800 observations (left and right, respectively). Labels are shortened concept questions generally of the form “Does the note mention the patient…?” Highlighted labels correspond to the true color-coded concepts in Section 4.2.
@@ 4.2 CBMs for classifying clinical notes with simulated outcomes
To objectively measure how well BC-LLM can recover true concepts, we conducted a simulation study where the true concepts are known. To assess BC-LLM in settings where the LLM has no prior knowledge about the target, we simulated $Y$ using a LR model with five Social Determinants of Health as inputs, which were annotated from real-world patient notes in MIMIC-IV [53, 54]: